x y
1 0
2 – 1
8 – 7
10 ?
20 ?
40 ?
Copy the equation. Replace b with a number that shows the relationship indicated in the table. Then copy the table and complete it using that equation.
y = – x + b

Answer by stanbon(75887)

x y
1 0
2 – 1
8 – 7
10 ?
20 ?
40 ?
Copy the equation. Replace b with a number that shows the relationship indicated in the table. Then copy the table and complete it using that equation.
y = – x + b
--------------------
Use the two points (1,0) and (2,1) to find the equation:
--------
slope = (1-0)/(2-1) = 1
---
y = mx+b
1 = 1*2+b
b = -1
-----
Equation:
y = x -1
----
x y
1 0
2 – 1
8 – 7
10 9
20 19
40 39
======
Cheers,
Stan H.